# The role of turbulence in setting the phase of the ISM and implications for the star formation rate

### v2

@@ 7.2 Heating/cooling model and threshold density
Figure 13: Theoretical temperature-density and pressure-density curve obtained from the cooling model of Wolfire et al. (2003), as described in Section 2.5, for different values of the UV field strength $G_{0}$. The gray line indicates the chosen temperature threshold for CNM.
We use the cooling/heating model from Wolfire et al. (2003) described in Section 2.5. The combination of processes leads to the curves shown in Figure 13. The only parameter we vary here is the UV field strength $G_{0}$. A strong UV field will boost photoelectric heating and move the WNM – CNM transition to higher densities, while leaving the shape of the curve relatively unchanged.
@@ 7.3 Density PDF generated by turbulence
It is well-established that turbulence generates log-normal-like density PDFs in isothermal media and that the Mach number controls the width of the distribution (Federrath et al., 2008, 2010a), as highly turbulent regions contain stronger density fluctuations. As a function of the density contrast $\delta=\ln(n/n_{0})$, where $n_{0}$ is the average number density of the region, the density PDF of a turbulent isothermal medium can be described by
@@ 7.7 Caveats and further improvements
A first avenue is to take into account the magnetic field. Molina et al. (2012) found that this can be achieved by including an additional factor in $S$:
@@
It has been pointed out that the PDF may not be exactly log-normal, but has a skewness to it especially at high Mach numbers (Hopkins, 2013; Hennebelle et al., 2024; Brucy et al., 2024). This may be important in the CNM. For the low Mach numbers that characterise the WNM, the log-normal form is a good approximation. We tried replacing the log-normal by this skewed version, dubbed the Castaing-Hopkins PDF (Castaing, 1996; Hopkins, 2013). While this produces a better approximation for the transition to LNM and an improved peak shape for the CNM, it introduced a sharp cut-off at high densities and essentially predicted zero gas with densities above 100 cm-3. One could argue that at this point gravity would significantly alter the density distribution, producing the additional power-law-like tail (Burkhart et al., 2017).
@@
That we do not capture the characteristic LNM minimum and CNM maximum, is likely due to the fact that we do not actually take into account that the LNM is an unstable equilibrium. The gas is more likely to be in one of the stable states, that is WNM or CNM. Furthermore, gas that is in the CNM and WNM has a preferred pressure and thus density. This pressure is given by the various vertical balancing forces in a galactic disk, among which feedback from star formation (Ostriker et al., 2010; Ostriker & Kim, 2022). Indeed, in Figure 5 we see the distribution of pressures is broad but peaked. This should somehow be taken into account.
@@ Acknowledgements
This research has received funding from the European Research Council synergy grant ECOGAL (Grant 855130). TC would like to thank all the members of the ECOGAL collaboration for discussions and comments which greatly improved this work. The authors acknowledge Interstellar Institute’s programs ”II4: The Grand Cascade”, as well as the Paris-Saclay University’s Institut Pascal for hosting discussions that triggered this work and nourished the development of the ideas behind it. This work was granted access to HPC resources of CINES and CCRT under the allocation x2020047023 made by GENCI (Grand Equipement National de Calcul Intensif). NB acknowledges support from the ANR BRIDGES grant (ANR-23-CE31-0005). In addition to ECOGAL, RSK and SCOG acknowledge funding from the German Excellence Strategy via the Heidelberg Excellence Cluster STRUCTURES (EXC-2181/1 - 390900948). TC acknowledges funding received from the European High Per- formance Computing Joint Undertaking (JU) and Belgium, Czech Republic, France, Germany, Greece, Italy, Norway, and Spain under grant agreement no. 101093441 (SPACE).

### v3

@@ 7.2 Heating/cooling model and threshold density
Figure 13: Theoretical temperature-density and pressure-density curve obtained from the cooling model of Wolfire_et_al2003, as described in Section 2.5, for different values of the UV field strength $G_{0}$. The gray line indicates the chosen temperature threshold for CNM.
We use the cooling/heating model from Wolfire_et_al2003 described in Section 2.5. The combination of processes leads to the curves shown in Figure 13. The only parameter we vary here is the UV field strength $G_{0}$. A strong UV field will boost photoelectric heating and move the WNM – CNM transition to higher densities, while leaving the shape of the curve relatively unchanged.
@@ 7.3 Density PDF generated by turbulence
It is well-established that turbulence generates log-normal-like density PDFs in isothermal media and that the Mach number controls the width of the distribution (Federrath_et_el2008; federrath10), as highly turbulent regions contain stronger density fluctuations. As a function of the density contrast $\delta=\ln(n/n_{0})$, where $n_{0}$ is the average number density of the region, the density PDF of a turbulent isothermal medium can be described by
@@ 7.7 Caveats and further improvements
A first avenue is to take into account the magnetic field. Molina_et_al2012 found that this can be achieved by including an additional factor in $S$:
@@
It has been pointed out that the PDF may not be exactly log-normal, but has a skewness to it especially at high Mach numbers (Hopkins2013; Hennebelle_et_al2024; Brucy_et_al2024). This may be important in the CNM. For the low Mach numbers that characterise the WNM, the log-normal form is a good approximation. We tried replacing the log-normal by this skewed version, dubbed the Castaing-Hopkins PDF (Castaing1996; Hopkins2013). While this produces a better approximation for the transition to LNM and an improved peak shape for the CNM, it introduced a sharp cut-off at high densities and essentially predicted zero gas with densities above 100 cm-3. One could argue that at this point gravity would significantly alter the density distribution, producing the additional power-law-like tail (Burkhart_et_al2017).
@@
That we do not capture the characteristic LNM minimum and CNM maximum, is likely due to the fact that we do not actually take into account that the LNM is an unstable equilibrium. The gas is more likely to be in one of the stable states, that is WNM or CNM. Furthermore, gas that is in the CNM and WNM has a preferred pressure and thus density. This pressure is given by the various vertical balancing forces in a galactic disk, among which feedback from star formation (Ostriker_et_al2010; Ostriker_Kim2022). Indeed, in Figure 5 we see the distribution of pressures is broad but peaked. This should somehow be taken into account.
@@ Acknowledgements
We thank the anonymous reviewers for their useful comments that led to significant improvements to this work. This research has received funding from the European Research Council synergy grant ECOGAL (Grant 855130). TC would like to thank all the members of the ECOGAL collaboration for discussions and comments which greatly improved this work. The authors acknowledge Interstellar Institute’s programs ”II4: The Grand Cascade”, as well as the Paris-Saclay University’s Institut Pascal for hosting discussions that triggered this work and nourished the development of the ideas behind it. This work was granted access to HPC resources of CINES and CCRT under the allocation x2020047023 made by GENCI (Grand Equipement National de Calcul Intensif). NB acknowledges support from the ANR BRIDGES grant (ANR-23-CE31-0005). In addition to ECOGAL, RSK and SCOG acknowledge funding from the German Excellence Strategy via the Heidelberg Excellence Cluster STRUCTURES (EXC-2181/1 - 390900948). TC acknowledges funding received from the European High Per- formance Computing Joint Undertaking (JU) and Belgium, Czech Republic, France, Germany, Greece, Italy, Norway, and Spain under grant agreement no. 101093441 (SPACE).